OpenStudy (anonymous):

Does this set of numbers form a Pythagorean triple? Explain please. :) 4,5,6

OpenStudy (anonymous):

no it does not. for it to form a pythagorean triangle it would have to conform with this rule - a^2+b^2=c^2. the longer side , in this case 6, is the hypotenuse, represented by c in the equation above. so it follows ; 4^2+5^2=41. 6^2 = 36 therefore 4^2 + 5^2 does not equal 6^2

OpenStudy (koikkara):

Hai Sweet Friend, @TheKylieeM WeLcoMe to OpeN stuDy.. !! Wondering If you are satisfied with this answer, Please Click "Best Response" and close this Question. Thank You ! Keep in touch with Open Study.. Bye !